OpenStudy (welshfella):

Integrate cot x csc^2x dx

OpenStudy (welshfella):

I'm trying the substitution u = cot x and using the identity csc^2 x = 1 + cot^2 x

OpenStudy (welshfella):

hold on its cot^x = 1 + csc^2 x right?

OpenStudy (anonymous):

\[\int\limits_{}^{}\cot x \csc ^2x dx=\int\limits_{}^{}\frac{ cosx }{ \sin^3x }dx\]

OpenStudy (welshfella):

and how would you proceed from there?

OpenStudy (amistre64):

if u = cot(x) what is du?

OpenStudy (anonymous):

u sub\[u=\sin x\]\[du=\cos x dx\]\[dx=\frac{ 1 }{ \cos x }\]

OpenStudy (welshfella):

du/dx ln sin x

OpenStudy (welshfella):

thats where i got stuck

OpenStudy (amistre64):

i recall D[tan(x)] = sec^2(x) D[cot(x)] = -csc^2(x)

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{ \cos x }{ u^3 }\frac{ 1 }{ \cos x }du\]

OpenStudy (amistre64):

\[\int cot(x)~csc^2(x)~dx\implies -\int u~du\]

OpenStudy (anonymous):

ya wanna take over am?

OpenStudy (welshfella):

ah - i was using ln sin x instead of -csc^2 x i can see it now

OpenStudy (amistre64):

take over? nah, just a perspective :)

OpenStudy (anonymous):

heh i have my own problems to solve too actually, not working too well

OpenStudy (amistre64):

welsh is a smart fella :)

OpenStudy (anonymous):

but i cant put it here cause it doesnt belong to maths

OpenStudy (welshfella):

yes I see it now thanks for your help

OpenStudy (welshfella):

I looked up the integral instead of the derivative!!!! that wasn't so smart!!!

OpenStudy (welshfella):

lol

OpenStudy (anonymous):

lol that happens alot

OpenStudy (welshfella):

ty both

OpenStudy (anonymous):

engineering department seems dead